OpenStudy (anonymous):

After making a $1,000 down payment, Mavis wants to finance the remaining $9,600 on the purchase price of a new car. She will barrow the money for five years at an annual interest rate of 10%. How much will she pay in interest on the loan.

OpenStudy (anonymous):

\left\{\left\{W\to \frac{P R (1+R)^N}{-1+(1+R)^N}\right\}\right\}